# Revisiting the Orbital Dynamics of the Hot Jupiter WASP–12b with New Transit Times

### v1

@@ Appendix A Graphical Representation of individual transit events taken from TESS, ETD and Exoclock
We have represented the TAP model fits for all 119 TESS light curves of WASP-12b in Figures 59. Similarly, 97 light curves from ETD and 34 light curves from Exoclock have been represented in Figure 1013 and Figure 1415.
Figure 5: The normalized relative flux of WASP-12b as a function of the time (the offset from mid-transit time and in TDB-based BJD) of individual transit observed by TESS between epochs (3965 - 4546) : the points are the data of raw flux, solid red lines are best-fit models for model flux, and E is the calculated epoch number.
@@
Figure 10: The normalized relative flux of WASP-12b as a function of the time (the offset from mid-transit time and in TDB-based BJD) of individual transits taken from ETD between epochs (602 - 1934) : here the points, solid red lines, and E represent the same elements as described in Figure 5.
@@
Figure 14: The normalized relative flux of WASP-12b as a function of the time (the offset from mid-transit time and in TDB-based BJD) of individual transits taken from Exoclock between epochs (974 - 4296) : here the points, solid red lines, and E represent the same elements as described in Figure 5.

### v2

@@ Appendix A Graphical Representation of individual transit events taken from TESS, ETD and Exoclock
We have represented model fits obtained from juliet for all 119 TESS light curves of WASP-12 b in Figures 37. Similarly, 97 light curves from ETD and 34 light curves from Exoclock have been represented in Figure 811 and Figure 1213.
Figure 3: The upper panel shows the relative normalized flux of WASP-12 b as a function of time (expressed as the offset from the mid-transit time in TDB-based BJD), together with the epoch number of individual transits observed by TESS between epochs 3965 and 4546. The flux error bars are shown in magenta. The solid curve is the best-fit light curve model obtained through juliet. The lower panel displays the corresponding residuals with their error bars.
@@
Figure 8: The upper panel shows the relative normalized flux of WASP-12 b as a function of time (expressed as the offset from the mid-transit time in TDB-based BJD), together with the epoch number of individual transits observed by ETD between epochs 602 and 1934. The flux error bars are shown in magenta and the solid curve is the best-fit light curve model obtained through juliet. The lower panel displays the corresponding residuals with their error bars.
@@
Figure 12: The upper panel shows the relative normalized flux of WASP-12 b as a function of time (expressed as the offset from the mid-transit time in TDB-based BJD), together with the epoch number of individual transits observed by Exoclock between epochs 974 and 4296. The flux error bars are shown in magenta and the solid curve is the best-fit light curve model obtained through juliet. The lower panel displays the corresponding residuals with their error bars.